OpenStudy (anonymous):

What do I do to solve this? "A plot of land is in the shape of a triangle. If one side is x meters, a second side is (2x-3) meters, and a third side is (3x-5) meters, express the perimeter of the lot as a simplified expression in x."

OpenStudy (anonymous):

for this one you can add up the lengths of each side: x + (2x - 3) + (3x - 5) try it out and see what you get

OpenStudy (anonymous):

well, I started trying to work it out, and it looks something like this 3x-4x+4x-5x but I'm not sure if that's right or if I'm doing something wrong :P

OpenStudy (anonymous):

or is it supposed to look like this 2x-3x+3x-5? x.x this is killin' me.

OpenStudy (anonymous):

All right let's break it down: x + (2x - 3) + (3x - 5) let's drop the parentheses to get: x + 2x - 3 + 3x - 5 first combine the x + 2x = 3x to obtain: 3x - 3 + 3x - 5 then combine the 3x + 3x = 6x to obtain: 6x - 3 - 5 for clarity, rewrite 6x - 3 - 5 as 6x + (-3) + (-5) this gives us: 6x + (-8) = 6x - 8

OpenStudy (anonymous):

k, let me try work that out on my own like you did it.

OpenStudy (anonymous):

Darn computer lost connection for a moment... was gonna ask a question about dropping the parenthesis. To drop the parenthesis, do you use (what I think is called) the distributive property? Where you multiply the stuff in the parenthesis by the number before it? If so, is there an unseen 1 in front of these parenthesis?

OpenStudy (anonymous):

Sorry if I ask a lot of questions. I tend to analyze EVERYTHING I come in contact with lol. My friends even tell me to stop analyzing things XD

OpenStudy (anonymous):

It complicates even the simplest tasks sometimes.

OpenStudy (anonymous):

btw, I did work the problem out how you did and got the same answer. so thank you.

OpenStudy (anonymous):

oh, so when parentheses, so long as there is a + sign before the parentheses, it's OK to drop them. for example: 1 + (3 - x) = 1 + 3 - x but if there is something in front of the parentheses like a number, then you'd have to distribute: 1 + 2(3 - x) = 1 + 6 - 2x likewise, if there is a minus sign in front of the parentheses, then you'd treat it as a + -1 like: 1 - (3 - x) = 1 + -1(3 - x) = 1 + -3 + x = 1 - 3 + x

OpenStudy (anonymous):

I gotta go. But thank you again for all your help.

OpenStudy (anonymous):

sure thing!